Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

die and coin

Expert replies
by maihuna » Fri Aug 14, 2009 9:08 am
Consider the experiment of throwing a die, if a multiple of 3 comes up, throw the die again and if any other number comes, toss a coin. Find the probability of the event ‘the coin shows a tail’, given that ‘at least one die shows a 3’.
Charged up again to beat the beast :)
Join the discussion
Source: — Problem Solving |

by DanaJ » Fri Aug 14, 2009 12:00 pm
:shock:

Probabilities... Hmm... Strange question, but interesting nonetheless.

From 1 to 6 you get two multiples of 3: 3 and 6. You will need to get 3 and 6 for your first toss. Then it gets a bit more complicated (or at least that's what I understand from if any other number comes). If you got a 6 the first time, now you need a 3 and vice versa. The coin toss will be simpler, since there's a fixed 50% chance of getting tails.

So two cases, each with the same probability. Getting a 3 the first time is 1/6, while getting a 6 the second time is again 1/6. Getting the tails will be 1/2. The probability for this event will be 1/6 * 1/6 * 1/2 = 1/72. Use the same reasoning to conclude that getting a 6 the first time and a 3 the second time yields the same probability.

Add them up to get 1/72 + 1/72 = 2/72 = 1/36.
Join the discussion

by kaulnikhil » Fri Aug 14, 2009 12:18 pm
s it 1/12??
Join the discussion

by maihuna » Fri Aug 14, 2009 10:37 pm
DanaJ wrote::shock:

Probabilities... Hmm... Strange question, but interesting nonetheless.

From 1 to 6 you get two multiples of 3: 3 and 6. You will need to get 3 and 6 for your first toss. Then it gets a bit more complicated (or at least that's what I understand from if any other number comes). If you got a 6 the first time, now you need a 3 and vice versa. The coin toss will be simpler, since there's a fixed 50% chance of getting tails.

So two cases, each with the same probability. Getting a 3 the first time is 1/6, while getting a 6 the second time is again 1/6. Getting the tails will be 1/2. The probability for this event will be 1/6 * 1/6 * 1/2 = 1/72. Use the same reasoning to conclude that getting a 6 the first time and a 3 the second time yields the same probability.

Add them up to get 1/72 + 1/72 = 2/72 = 1/36.
Answer is 0.
Charged up again to beat the beast :)
Join the discussion

by DanaJ » Fri Aug 14, 2009 11:17 pm
I'd like to see the official explanation, if the answer is 0. Maybe I misunderstood the problem.
Join the discussion

by [email protected] » Mon Aug 17, 2009 1:11 am
Is this a GMAT kind of question?
Join the discussion